Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

three-story apartment building

Expert replies
by vikram4689 » Sun Aug 19, 2012 4:09 am
In a certain three-story apartment building, there are 11 tenants on the first floor, 3 tenants on the second floor, and 7 tenants on the third floor. If two tenants are selected at random to be co-chairs of the tenants' association, what is the probability that the two co-chairs will be from adjacent floors?


(A) 7/60

(B) 9/70

(C) 11/60

(D) 11/70

(E) 9/35

how can it be solved using INDIVIDUAL PROBABILITY. i mean don't solve using desired pairs/total pairs
Premise: If you like my post
Conclusion : Press the Thanks Button ;)
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Sun Aug 19, 2012 6:31 am
vikram4689 wrote:In a certain three-story apartment building, there are 11 tenants on the first floor, 3 tenants on the second floor, and 7 tenants on the third floor. If two tenants are selected at random to be co-chairs of the tenants' association, what is the probability that the two co-chairs will be from adjacent floors?
(A) 7/60
(B) 9/70
(C) 11/60
(D) 11/70
(E) 9/35
Always ask, "What needs to happen for the desired outcome to occur?"
We get: P(co-chairs from adjacent floors) = P(one from 1st floor then one from 2nd floor OR one from 2nd floor then one from 3rd floor OR one from 2nd floor then one from 1st floor OR one from 3rd floor then one from 2nd floor).
= (11/21)(3/20) + (3/21)(7/20) + (3/21)(11/20) + (7/21)(3/20)
= 33/420 + 21/420 + 33/420 + 21/420
= 108/420
= 9/35
= E

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Brent@GMATPrepNow » Sun Aug 19, 2012 7:28 am
vikram4689 wrote:In a certain three-story apartment building, there are 11 tenants on the first floor, 3 tenants on the second floor, and 7 tenants on the third floor. If two tenants are selected at random to be co-chairs of the tenants' association, what is the probability that the two co-chairs will be from adjacent floors?
(A) 7/60
(B) 9/70
(C) 11/60
(D) 11/70
(E) 9/35
We can also solve this question using counting techniques (as is the case with many probability questions).
In this case, it might be useful to use the complement. That is P(co-chairs from adjacent floors) = 1 - P(co-chairs not from adjacent floors)

P(co-chairs not from adjacent floors) = (# of outcomes where selections are not adjacent)/(total # outcomes)

As always, calculate the denominator first.
total # outcomes = total number of ways to select 2 tenants from 21 tenants
This can be accomplished in 21C2 ways (this is combination notation)


# of outcomes where selections are not adjacent
There are two ways this can happen:
1) The two people are selected from floor #1 and/or floor #3
2) The two people are selected from floor #2

1) The two people are selected from floor #1 and/or floor #3
To determine the number of possible outcomes, take the 18 tenants from floors 1 and 3, and select 2 of them.
This can be accomplished in 18C2 ways

2) The two people are selected from floor #2
To determine the number of possible outcomes, take the 3 tenants from floor #2, and select 2 of them.
This can be accomplished in 3C2 ways


So, P(co-chairs not from adjacent floors) = (18C2 + 3C2)/(21C2)
= (153 + 3)/(210)
= 156/210
= 26/35


And finally, P(co-chairs from adjacent floors) = 1 - P(co-chairs not from adjacent floors)
= 1 - 26/35
= 9/35
= E

Cheers,
Brent

If anyone is interested, we have a free video on calculating combinations in your head: https://www.gmatprepnow.com/module/gmat-counting?id=789
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion